# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v2

@@ 9.4. Irrelevant terms
We have $\operatorname{deg}^{\mathrm{Dr}}E_{0}=-\sum_{j=1}^{N}\alpha_{j}$ and $\operatorname{deg}^{\mathrm{Dr}}F_{0}=\sum_{j=1}^{N}\alpha_{j}$.
@@
Let us write $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ as a sum of $Q$-homogeneous terms. We say that a term is irrelevant if its degree is not a multiple of $\alpha_{i}$.
@@
Every irrelevant term of $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ lies in $U_{d_{i}=1,+}$ ($1\leq i\leq N$). More specifically, it can be written as a sum of monomials consisting of letters $F_{1},\cdots,$ $F_{m-1},\tilde{E}_{m},\cdots\tilde{E}_{i-1},\tilde{E}_{i+1},\cdots,\tilde{E}_{n},F_{n+1},\cdots,F_{N},F_{0}$, such that: (a) $1\leq m\leq i$ and $i\leq n\leq N$, (b) either $m<i$ or $n>i$, (c) each letter occurs exactly once.
@@
In analogy to (9.9), there exists a polynomial $P(\bullet)$ such that
$\mathbf{T}_{1}\cdots\mathbf{T}_{i-1}(B_{i})=P(B_{2},\cdots,B_{i})B_{1}-qB_{1}P(B_{2},\cdots,B_{i}).$
@@
for $2\leq k\leq B_{i}$. Therefore, Lemma 8.4 implies that
(9.16) $\pi^{-i}\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=$ $\ [P(B_{N-i+2},\cdots,B_{N}),\mathbf{T}_{N-i+1}\cdots\mathbf{T}_{2}(B_{1})]_{q}.$
By Lemma 9.2, $\eta(\mathbf{T}_{N-i+1}\cdots\mathbf{T}_{2}(B_{1}))$ can be written as a sum of monomials consisting of letters $\tilde{E}_{1},\cdots,\tilde{E}_{l},F_{l+1},\cdots,F_{N-i+1}$, and, by Lemma 9.3, $\eta(P(B_{N-i+2},\cdots,B_{N}))$ as a sum of monomials consisting of letters $F_{N-i+2},\cdots,F_{k-1},\tilde{E}_{k},\cdots,\tilde{E}_{N}$ (each letter occurring exactly once).
Next, we claim that $\eta(\pi^{-i}\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ contains no monomial with letters $\tilde{E}_{N-i+1}$ and $F_{N-i+2}$. This implies that the only monomials containing the letter $\tilde{E}_{N-i+1}$ are those not containing any letters $F_{k}$. Also note that none of the monomials in $\eta(\pi^{-i}\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ contain the letters $\tilde{E}_{0}$ or $F_{0}$. Therefore, applying $\pi^{i}$, we deduce that none of the monomials in $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ contain $\tilde{E}_{i}$ or $F_{i}$. Moreover, the monomials in $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ can be divided into the following kinds:
  • consisting of letters $\tilde{E}_{j}$ ($i\neq j\in\mathbb{I}$), each occurring exactly once,
  • consisting of letters $F_{j}$ ($i\neq j\in\mathbb{I}$), each occurring exactly once,
  • consisting of the letter $F_{0}$, as well as letters $F_{1},\cdots,F_{m-1},\tilde{E}_{m},\cdots\tilde{E}_{i-1},\tilde{E}_{i+1},\cdots,$ $\tilde{E}_{n},F_{n+1},\cdots,F_{N}$, and satisfying the conditions in the statement of the proposition.
Using Lemma 9.4, we conclude that precisely the monomials of the third kind are irrelevant and that they lie in $U_{d_{i}=1,+}$.
It remains to prove the claim. Suppose that $\eta(\pi^{-i}\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ contained monomials with letters $\tilde{E}_{N-i+1}$ and $F_{N-i+2}$. By (9.16), such monomials arise from taking the $q$-commutator of monomials in $P(B_{N-i+2},\cdots,B_{N})$ consisting of letters $F_{N-i+2},\cdots,F_{k-1}$, $\tilde{E}_{k},\cdots,\tilde{E}_{N}$ (with $k\geq N-i+3$), and monomials in $\mathbf{T}_{N-i+1}\cdots\mathbf{T}_{2}(B_{1})$ consisting of letters $\tilde{E}_{1},\cdots,\tilde{E}_{N-i+1}$. Let us take two such monomials and write them as $aF_{N-i+2}b$ and $c\tilde{E}_{N-i+1}d$, respectively, where $a,b$ are monomials in $F_{N-i+3},\cdots,F_{k-1},\tilde{E}_{k},\cdots,\tilde{E}_{N}$ (with $k\geq N-i+3$), and $c,d$ are monomials in $\tilde{E}_{1},\cdots,\tilde{E}_{N-i}$, Since $a,b$ commute with $c\tilde{E}_{N-i+1}d$, and $c,d$ commute with $aF_{N-i+2}b$, we get
(9.17) $[aF_{N-i+2}b,c\tilde{E}_{N-i+1}d]_{q}=ac[F_{N-i+2},\tilde{E}_{N-i+1}]_{q}bd=0,$
since $[F_{N-i+2},\tilde{E}_{N-i+1}]_{q}=0$, as in (9.8). ∎
9.5. Factorization theorem in higher rank
In this section we generalize the factorization theorem to higher ranks. We need the following lemma.
Proposition 9.6.
We have
(9.18) $\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))$$\equiv$ $\ T_{\omega_{i}^{\prime}}(\eta(B_{i}))$ $\mod U_{d_{i}=1,+},$
for each $1\leq i\leq N$.
Proof.
By Proposition 9.5, all the irrelevant terms of $\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))$ lie in $U_{d_{i}=1,+}$. Therefore, it suffices to check that the relevant terms in $\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))$ and $T_{\omega_{i}^{\prime}}(\eta(B_{i}))$ coincide. The proof is analogous to the first stage of the proof of Proposition 9.5. More precisely, using the definitions of the two braid group actions, Lemma 9.1 and induction, one shows that the term of $\eta(\mathbf{T}_{1}\cdots\mathbf{T}_{i-1}(B_{i}))$ consisting only of letters $F_{l}$ (resp. $\tilde{E}_{l}$) coincides with $T_{1}\cdots T_{i-1}(F_{i})$ (resp. $T_{1}\cdots T_{i-1}(\tilde{E}_{i})$. Using (9.16), we deduce that the term of degree $\alpha_{i}$ in
$\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))=\eta(\pi^{i}[P(B_{N-i+2},\cdots,B_{N}),\mathbf{T}_{N-i+1}\cdots\mathbf{T}_{2}(B_{1})]_{q})$
equals
$\mathbf{T}_{\omega^{\prime}_{i}}(F_{i})=\pi^{i}[P(F_{N-i+2},\cdots,F_{N}),T_{N-i+1}\cdots T_{2}(F_{1})]_{q}.$
An analogous argument holds for the terms of degree $-\alpha_{i}$, with $F_{l}$ replaced by $\tilde{E}_{l}$. ∎
We now derive the main result of this section, the following compatibility theorem.
Theorem 9.7 (Compatibility theorem).
@@ 9.5. Factorization theorem in higher rank
Thanks to Theorem 9.7, we can apply our rank one results from §4-7 in the general case.

### v3

@@ 9.4. Weak compatibility
We have $\operatorname{deg}^{\mathrm{Dr}}E_{0}=-\theta=-\sum_{j=1}^{N}\alpha_{j}$ and $\operatorname{deg}^{\mathrm{Dr}}F_{0}=\theta=\sum_{j=1}^{N}\alpha_{j}$.
@@
Let us write $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ as a sum of $Q$-homogeneous terms. We say that a term is irrelevant if its degree is not ±$\pm\alpha_{i}$. The proposition below is our first result on the compatibility of the two braid group actions.
@@
Let $1\leq i\leq N$. Every irrelevant term of $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ lies in $U_{d_{i}=1,+}$. Hence
(9.15) $\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))$$\equiv$ $\ T_{\omega_{i}^{\prime}}(\eta(B_{i}))$ $\mod U_{d_{i}=1,+}.$
@@
By Lemma 9.4,
$\mathbf{T}_{1}\cdots\mathbf{T}_{i-1}(B_{i})=[P_{i-1}(B_{i},\cdots,B_{2}),B_{1}]_{q}.$
@@
for $2\leq k\leq B_{i}$. Therefore, Lemmas 8.4 and 9.3-9.4 imply that
(9.18) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=$ $\ \pi^{i}[P_{i-1}(B_{N},\cdots,B_{N-i+2}),\mathbf{T}_{N-i+1}\cdots\mathbf{T}_{2}(B_{1})]_{q}$
(9.19) $=$ $\ \pi^{i}[P_{i-1}(B_{N},\cdots,B_{N-i+2}),P_{N-i+1}(B_{1},\cdots,B_{N-i+1})]_{q}$
(9.20) $=$ $\ \pi^{i}\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=N-i+2,...,N\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon_{l}\geq\varepsilon_{l+1},\\ l=1,...,N-i\end{subarray}}[P_{i-1}(\tilde{e}_{N}^{\varepsilon_{N}},\cdots,\tilde{e}_{N-i+2}^{\varepsilon_{N-i+2}}),P_{N-i+1}(\tilde{e}_{1}^{\varepsilon_{1}},\cdots,\tilde{e}_{N-i+1}^{\varepsilon_{N-i+1}})]_{q}$
(9.21) $=$ $\ \sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=1,...,i-1\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon_{l}\geq\varepsilon_{l+1}\geq\varepsilon_{0},\\ l=i+1,...,N-1\end{subarray}}[P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q}.$
By Lemma 9.5, each irrelevant term inside of the summation in (9.21) is already in $U_{d_{i}=1,+}$ unless $\varepsilon_{0}=+$. So suppose that $\varepsilon_{0}=+$ and that at least one of the superscripts $\varepsilon_{k}=-$. This forces $\varepsilon_{1}=-$ and $\varepsilon_{i+1}=\cdots=\varepsilon_{0}=+$. We then have
(9.23) $[P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),$ $\ P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q}=$
(9.24) $=$ $\ [[P_{i-2}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}}),F_{1}]_{q},[P_{N-i}(\tilde{E}_{i+1},\cdots,\tilde{E}_{N}),\tilde{E}_{0}]_{q}]_{q}$
(9.25) $=$ $\ [P_{N-i}(\tilde{E}_{i+1},\cdots,\tilde{E}_{N}),[P_{i-2}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}}),[F_{1},\tilde{E}_{0}]_{q}]_{q}]_{q}$
[…]
since, by (9.9), $[F_{1},\tilde{E}_{0}]_{q}=0$. We conclude that every irrelevant term in (9.21) is indeed in $U_{d_{i}=1,+}$. Moreover, the relevant terms are precisely
(9.27) $[P_{i-1}(\tilde{E}_{i-1},\cdots,\tilde{E}_{1}),P_{N-i+1}(\tilde{E}_{i+1},\cdots,\tilde{E}_{0})]_{q}=$ $\ T_{\omega_{i}^{\prime}}(\tilde{E}_{i}),$
(9.28) $[P_{i-1}({F}_{i-1},\cdots,{F}_{1}),P_{N-i+1}({F}_{i+1},\cdots,{F}_{0})]_{q}=$ $\ T_{\omega_{i}^{\prime}}({F}_{i}).$
This concludes the proof. ∎
Corollary 9.7 (Compatibility theorem).
@@
9.5. Strong compatibility
The goal of this section is to strengthen the results of the preceding subsection to pave the way for a higher rank generalization of the factorization theorem. Set $\mathfrak{C}_{i}=\mathfrak{C}^{-1}\mathbb{K}_{i}$. It is easily verified that
$\mathbf{T}_{\omega^{\prime}_{i}}(\mathbb{K}_{i})=\mathfrak{C}_{i}^{-1},\qquad T_{\omega^{\prime}_{i}}(K_{i})=K_{i}^{-1}.$
In the calculations below, we will frequently make use of a remainder term $Q_{i}$. The statement of the following lemma contains its definition.
Lemma 9.8.
For each $i\in\mathbb{I}_{0}$:
$\eta(A_{i,-1})=x^{-}_{i,1}-\mathfrak{C}_{i}x^{+}_{i,-1}K_{i}+Q_{i},\qquad Q_{i}\in U_{d_{i}\geq 1,+}.$
Hence
$\eta(A_{i,-1})\equiv\iota_{i}\eta(A_{-1})\quad\mod U_{d_{i}\geq 1,+}.$
Proof.
By [Bec94, Proposition 4.7] [^4] More precisely, in the first equality below, an adjustment by a sign function $o(i)$ which alternates on adjacent vertices is necessary. An analogous adjustment occurs in the formula for $A_{i,-1}$. Since we work with a fixed vertex $i\in\mathbb{I}_{0}$ most of the time, we may assume for simplicity that $o(i)=1$.,
(9.29) $x_{i,-1}^{+}=$ $\ T_{\omega_{i}}(E_{i})=T_{\omega^{\prime}_{i}}(-F_{i}K_{i})=-T_{\omega^{\prime}_{i}}(F_{i})K_{i}^{-1},$
(9.30) $x_{i,1}^{-}=$ $\ T_{\omega_{i}}(F_{i})=T_{\omega^{\prime}_{i}}(-K_{i}^{-1}E_{i})=-K_{i}T_{\omega^{\prime}_{i}}(E_{i}).$
Moreover, by [LW21, Theorem 3.13],
$A_{i,-1}=\mathbf{T}_{\omega_{i}}(B_{i})=\mathbf{T}_{\omega^{\prime}_{i}}(\mathbb{K}_{i}^{-1}B_{i})=\mathfrak{C}_{i}\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}).$
Hence, by Proposition 9.6 and (9.29)–(9.30),
$\eta(A_{i,-1})=\mathfrak{C}_{i}T_{\omega^{\prime}_{i}}(F_{i}-K_{i}^{-1}\mathbb{K}_{i}E_{i})+Q_{i}=-\mathfrak{C}_{i}x^{+}_{i,-1}K_{i}+x^{-}_{i,1}+Q_{i},$
completing the proof. ∎
Lemma 9.9.
We have
$\eta(H_{i,1})\equiv\iota_{i}\eta(H_{1})+q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}}\quad\mod U_{d_{i}\geq 2,+}.$
Proof.
We have
(9.31) $\eta(H_{i,1})=$ $\ q^{2}\mathfrak{C}_{i}^{-1}\eta([A_{i,-1},A_{i,0}]_{q_{i}^{-2}})$
(9.32) $=$ $\ q^{2}\mathfrak{C}_{i}^{-1}[\iota_{i}\eta(A_{-1})+Q_{i},\iota_{i}\eta(A_{0})]_{q^{-2}}$
(9.33) $=$ $\ \iota_{i}\eta(H_{1})+q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}+\tilde{E}_{i}]_{q^{-2}}.$
Since $[Q_{i},\tilde{E}_{i}]_{q^{-2}}\in U_{d_{i}\geq 2,+}$, the lemma follows. ∎
Proposition 9.10.
For all $r\in{\mathbb{Z}}$:
$\eta(A_{i,r})\equiv\iota_{i}\eta(A_{r})\quad\mod U_{d_{i}\geq 1,+}.$
Proof.
The result holds for $r=-1,0$ by Proposition 9.6. The general case is proven using upward and downward induction. We explicitly do the former, the latter being analogous. Relation (8.36) implies that
(9.34) $A_{i,r+1}=[2]_{q}^{-1}[H_{i,1},A_{i,r}]+\mathfrak{C}A_{i,r-1}.$
Using Lemma 9.9, let us write $\eta(H_{i,1})=\iota_{i}\eta(H_{1})+q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}}+R$, with $R\in U_{d_{i}\geq 2,+}$. By induction, we may assume that
$\eta(A_{i,r-1})\equiv\iota_{i}\eta(A_{r-1}),\quad[\iota_{i}\eta(H_{1}),\eta(A_{i,r})]\equiv\iota_{i}\eta([H_{1},A_{r}])$
modulo $U_{d_{i}\geq 1,+}$. Therefore, it suffices to show that
$[q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}}+R,\eta(A_{i,r})]\in U_{d_{i}\geq 1,+}.$
By induction, we can write $\eta(A_{i,r})=\iota_{i}\eta(A_{r})+S$, with $S\in U_{d_{i}\geq 1,+}$. On the other hand, by Lemma 4.5 and (8.45), $\iota_{i}\eta(A_{r})=x_{i,-r}^{-}+S^{\prime}$ with $S^{\prime}\in U_{[i],+}$, where $U_{[i],+}$ is the subalgebra of $U_{[i]}$ spanned by elements of positive degree (with respect to $\alpha_{i}$). We observe immediately that
$[R,\eta(A_{i,r})]+[q^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q^{-2}},S+S^{\prime}]\in U_{d_{i}\geq 1,+}.$
Therefore, it suffices to show that
(9.35) $[[Q_{i},F_{i}]_{q^{-2}},x_{i,-r}^{-}]\in U_{d_{i}\geq 1,+}.$
In fact, we show that (9.35) vanishes. This will be achieved in §9.6 below. ∎
9.6. Completion of the proof
We first state some auxiliary lemmas. Define the non-commutative polynomial $P^{\prime}_{k}(y_{1},\cdots,y_{k})$ over ${\mathbb{C}}$ by induction in the following way:
(9.36) $P^{\prime}_{1}(y_{1})=$ $\ y_{1},\qquad$ $P^{\prime}_{k+1}(y_{1},\cdots,y_{k})=$ $\ P^{\prime}_{k}(y_{1},\cdots,y_{k-1},[y_{k},y_{k+1}]_{q}),$
We have
(9.37) $P^{\prime}_{k}(y_{1},\cdots,y_{k})=$ $\ P^{\prime}_{l+1}(y_{1},\cdots,y_{l},P^{\prime}_{k-l}(y_{l+1},\cdots,y_{k})),$
for any $11\leq l\leq k-$ 1.
We say that a tuple $(y_{1},\cdots,y_{k})$ is almost commuting if $y_{m}y_{n}=y_{n}y_{m}$, for $|m-n|>1$.
Lemma 9.11.
The polynomials $P_{k}$ and $P^{\prime}_{k}$ are equal if $(y_{1},\cdots,y_{k})$ is almost commuting.
Proof.
The first statement is proven by induction. By definition, $P_{1}=P^{\prime}_{1}$ and $P_{2}=P^{\prime}_{2}$. Moreover,
(9.39) $P^{\prime}_{k+1}(y_{1},\cdots,y_{k+1})=$ $\ P^{\prime}_{2}(y_{1},P^{\prime}_{k}(y_{2},\cdots,y_{k+1}))$
(9.40) $=$ $\ P^{\prime}_{2}(y_{1},P_{k}(y_{2},\cdots,y_{k+1}))$
(9.41) $=$ $\ P_{k}(P^{\prime}_{2}(y_{1},y_{2}),y_{3},\cdots,y_{k+1})$
(9.42) $=$ $\ P_{k}(P_{2}(y_{1},y_{2}),y_{3},\cdots,y_{k+1})=P^{\prime}_{k+1}(y_{1},\cdots,y_{k+1}).\qed$
Lemma 9.12.
Let $12\leq i\leq n-$ 1. Then
  1. (1)
    $[\tilde{E}_{i},[\tilde{E}_{i-1},Y]_{q}]_{q}=q[E_{i},[\tilde{E}_{i-1}K_{i}^{-1},Y]_{q}]_{q^{2}}$ if $[K_{i}^{-1},Y]=0$,
  2. (2)
    $[[\tilde{E}_{i-1},Y]_{q},F_{i}]_{q^{-2}}=[\tilde{E}_{i-1},[Y,F_{i}]_{q^{-1}}]$,
  3. (3)
    $[[\tilde{E}_{i+1},Y]_{q},F_{i}]_{q^{-1}}=[\tilde{E}_{i+1},[Y,F_{i}]]$,
for any $Y\in U_{q}^{\mathbb{K}}L\mathfrak{g}$. If $[Y,F_{i}]=0$ then
(9.43) $[[\tilde{E}_{i-1},[\tilde{E}_{i+1},Y]_{q}]_{q},F_{i}]_{q^{-2}}=0.$
Proof.
The first three statements follow by a direct calculation. The fourth statement follows from the second and third if we substitute $[\tilde{E}_{i+1},Y]_{q}$ for the $Y$ in (2). ∎
Lemma 9.13.
Let $i-1\leq k\leq 1$. The element
$Z=[[P_{i-1}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},F_{k-1},\cdots,F_{1}),P_{N-i+1}(F_{i+1},\cdots,F_{N},F_{0})]_{q},F_{i}]_{q^{-2}}$
can be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$, with coefficients in ${\mathbb{C}}$. Hence
(9.44) $[Z,x_{i,s}^{-}]=0$
for all $s\in{\mathbb{Z}}$.
Similarly, if $i+1\leq k\leq n$ or $k=0$, then
$Z^{\prime}=[[P_{i-1}(F_{i-1},\cdots,F_{1}),P_{N-i+1}(\tilde{E}_{i+1},\cdots,\tilde{E}_{k},F_{k-1},\cdots F_{N},F_{0})]_{q},F_{i}]_{q^{-2}}$
can also be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$, with coefficients in ${\mathbb{C}}$. Hence
(9.45) $[Z^{\prime},x_{i,s}^{-}]=0.$
Proof.
We will only prove the statement for $Z$, the one for $Z^{\prime}$ being analogous. The equality (9.44) follows directly from relation (8.13) and $Z$ being a polynomial in $x_{j,r}^{+}$ $(j\neq i)$. Therefore, it suffices to prove the latter. Using (9.37) and Lemma 9.11, we have
(9.46) $[P_{i-1}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},$ $\ F_{k-1},\cdots,F_{1}),P_{N-i+1}(F_{i+1},\cdots,F_{N},F_{0})]_{q}=$
(9.47) $=$ $\ P^{\prime}_{i}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},F_{k-1},\cdots,F_{1},P^{\prime}_{N-i+1}(F_{i+1},\cdots,F_{N},F_{0}))$
(9.48) $=$ $\ P^{\prime}_{N}(\tilde{E}_{i-1},\cdots,\tilde{E}_{k},F_{k-1},\cdots,F_{1},F_{i+1},\cdots,F_{N},F_{0}).$
In the following, let us abbreviate
$\vec{F}$ ” $=$$F_{k{\shortminus}1},\cdots,F_{1},F_{i},F_{i+1},\cdots,F_{N},F_{0}$”.
Direct calculation using Lemma 9.12 shows that
(9.49) $Z=$ $\ {\shortminus}q[\tilde{E}_{i{\shortminus}1},P^{\prime}_{N}(F_{i},\tilde{E}_{i{\shortminus}2},\cdots,\tilde{E}_{k},F_{k{\shortminus}1},\cdots,F_{1},F_{i+1},\cdots,F_{N},F_{0})]$
(9.50) $=$ $\ {\shortminus}q[\tilde{E}_{i{\shortminus}1},P^{\prime}_{N}(\tilde{E}_{i{\shortminus}2},\cdots,\tilde{E}_{k},F_{k{\shortminus}1},\cdots,F_{1},F_{i},F_{i+1},\cdots,F_{N},F_{0})]$
(9.51) $=$ $\ {\shortminus}q[\tilde{E}_{i{\shortminus}1},[E_{i{\shortminus}2},P^{\prime}_{N{\shortminus}1}(E_{i{\shortminus}3},\cdots,E_{k}K_{i{\shortminus}2}^{-1}\cdots K_{k}^{-1},\vec{F})]]$
(9.52) $=$ $\ {\shortminus}q^{2}[E_{i{\shortminus}1},[E_{i{\shortminus}2}K_{i{\shortminus}1}^{-1},P^{\prime}_{N{\shortminus}1}(E_{i{\shortminus}3},\cdots,E_{k}K_{i{\shortminus}2}^{-1}\cdots K_{k}^{-1},\vec{F})]_{q^{-1}}]_{q^{2}}$
(9.53) $=$ $\ {\shortminus}q^{2}[E_{i{\shortminus}1},[E_{i{\shortminus}2},[E_{i{\shortminus}3}K_{i{\shortminus}1}^{-1},P^{\prime}_{N{\shortminus}2}(E_{i{\shortminus}4},\cdots,E_{k}K_{i{\shortminus}2}^{-1}\cdots K_{k}^{-1},\vec{F})]]_{q^{-1}}]_{q^{2}}$
(9.54) $=$ $\ {\shortminus}q^{2}[E_{i{\shortminus}1},[E_{i{\shortminus}2},P^{\prime}_{N{\shortminus}1}(E_{i{\shortminus}3},E_{i{\shortminus}4},\cdots,E_{k}K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1},\vec{F})]_{q^{-1}}]_{q^{2}}$
(9.55) $=$ $\ q^{3}(q{\shortminus}q^{-1})[E_{i{\shortminus}1},[E_{i{\shortminus}2},P^{\prime}_{i{\shortminus}k{\shortminus}1}(E_{i{\shortminus}3},E_{i{\shortminus}4},\cdots,$
(9.56) $\ E_{k+1},E_{k}P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1})]_{q^{-1}}]_{q^{2}}.$
Therefore, it suffices to show that $P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1}$ can be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$. First assume that $i>1$. We claim that
$P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1}=P^{\prime}_{i-k}(x_{i-1,-1}^{+},E_{i-2},\cdots,E_{k})=:Y.$
Let
$\zeta=\pi^{i-1}[N-i+2,N]\cdots[2,i].$
Indeed,
(9.57) $Y=$ $\ P^{\prime}_{i-k}(x_{i-1,-1}^{+},E_{i-2},\cdots,E_{k})=T_{\omega_{i-1}}P^{\prime}_{i-k}(E_{i-1},E_{i-2},\cdots,E_{k})$
(9.58) $=$ $\ T_{\omega_{i-1}}T_{k}\cdots T_{i-2}(E_{i-1})=T_{\zeta}T_{[1,k]}(E_{k})$
(9.59) $=$ $\ -T_{\zeta}T_{[1,k-1]}(F_{k}K_{k})=-T_{\zeta}(P^{\prime}_{k}(F_{k},\cdots,F_{1})K_{1}\cdots K_{k})$
(9.60) $=$ $\ -T_{\pi^{i-1}}(P^{\prime}_{k}(F_{N+k-i+1},\cdots,F_{N-i+3},F_{1},\cdots,F_{N-i+2})\cdot$
(9.61) $\ \cdot K_{N+k-i+1}\cdots K_{N-i+3}K_{1}\cdots K_{N-i+2})$
(9.62) $=$ $\ -P^{\prime}_{k}(F_{k-1},\cdots,F_{1},F_{i},\cdots,F_{N},F_{0})K_{k-1}\cdots K_{1}K_{i}\cdots K_{N}K_{0}$
(9.63) $=$ $\ -P^{\prime}_{N{\shortminus}i+k+2}(\vec{F})K_{i{\shortminus}1}^{-1}\cdots K_{k}^{-1},$
proving the claim. The case $i=1$ is similar but easier, so we leave it to the reader. ∎
Proposition 9.14.
We have
$[[Q_{i},F_{i}]_{q^{-2}},x^{-}_{i,r}]=0.$
Proof.
By the definition of $Q_{i}$ and (9.21), we have
(9.64) $Q_{i}=\mathfrak{C}_{i}\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ k=1,...,i-1\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon_{l}\geq\varepsilon_{l+1}\geq\varepsilon_{0},\\ l=i+1,...,N-1\end{subarray}}[P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}}),P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q},$
where $(\varepsilon_{i-1},\varepsilon_{i+1})\in\{(+,-),(-,+),(+,+)\}$. Let us denote each summand on the RHS of (9.64), corresponding to a sequence $\vec{\varepsilon}=(\varepsilon_{i-1},\cdots\varepsilon_{1},\varepsilon_{i+1},\cdots\varepsilon_{N},\varepsilon_{0})$, by $P_{\vec{\varepsilon}}$.
It follows from Lemma 9.11 and (9.37) that
(9.65) $P_{i-1}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}})=$ $\ [\tilde{e}_{i-1}^{\varepsilon_{i-1}},P_{i-2}(\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}})]_{q},$
(9.66) $P_{N-i+1}(\tilde{e}_{i+1}^{\varepsilon_{i+1}},\tilde{e}_{i+2}^{\varepsilon_{i+2}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})=$ $\ [\tilde{e}_{i+1}^{\varepsilon_{i+1}},P_{N-i}(\tilde{e}_{i+2}^{\varepsilon_{i+2}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})]_{q}.$
Let us abbreviate $Y_{1}=P_{i-2}(\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}})$ and $Y_{2}=P_{N-i}(\tilde{e}_{i+2}^{\varepsilon_{i+2}},\cdots,\tilde{e}_{0}^{\varepsilon_{0}})$. Then
(9.67) $P_{\vec{\varepsilon}}=$ $\ [[\tilde{e}_{i-1}^{\varepsilon_{i-1}},Y_{1}]_{q},[\tilde{e}_{i+1}^{\varepsilon_{i+1}},Y_{2}]_{q}]_{q}$
(9.68) $=$ $\ (1-q)[\tilde{e}_{i-1}^{\varepsilon_{i-1}},[\tilde{e}_{i+1}^{\varepsilon_{i+1}},Y_{1}Y_{2}]_{q}]_{q}.$
If $(\varepsilon_{i-1},\varepsilon_{i+1})=(+,+)$, then (9.43) in Lemma 9.12 now implies that $[P_{\vec{\varepsilon}},F_{i}]_{q^{-2}}=0$. Otherwise, $[[P_{\vec{\varepsilon}},F_{i}]_{q^{-2}},x^{-}_{i,r}]=0$ follows directly from (9.44)–(9.45) in Lemma 9.13. ∎
9.7. Factorization theorem in higher rank
We can now generalize the factorization and coproduct theorems to higher ranks.